write equation that has a slope of -2/3 and contains (-6,-4).

Answer by mananth(16946)

m=slope=-2/3
point(-6,-4)
equation in slope point form
(y-y1)=m(x-x1) Where (x1,y1) are the coordinates and m is the slope
substitute
y-(-4)=(-2/3)(x-(-6))
y+4 = (-2/3)x -4
multiply equation by 3
3y+12= -2x-12
3y=-12-2x-12
3y=-2x-24
y=(-2/3)x -8